OpenStudy (anonymous):

(Pls help me pretty pretty please) (the question and choices is at the bottom. I'll give you a medal D: )

OpenStudy (anonymous):

I can't do this. I just gotta go work at a hotdog stand

OpenStudy (anonymous):

>_____________________< Imma cry

OpenStudy (anonymous):

I see you here again. Hiii D:

OpenStudy (anonymous):

I told you just to have someone help you with the system of equations.

OpenStudy (anonymous):

I did D:

OpenStudy (anonymous):

And what did ya get?

OpenStudy (anonymous):

Nothing! >.< no one helped me

OpenStudy (anonymous):

*curls up in a ball*

OpenStudy (anonymous):

I'll help

OpenStudy (anonymous):

just a sec I need to solve this

OpenStudy (anonymous):

Go for it, mystical flames man

OpenStudy (anonymous):

You will help me with the system of equation? yayyy

OpenStudy (anonymous):

\(^_^)/

OpenStudy (anonymous):

okay, so

OpenStudy (anonymous):

mSRT + mRST = 8x (exterior angle is sum of interior angles)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay mSRT + (7x – 54) = 8x

OpenStudy (anonymous):

take it to the other side, and mSRT = 8x – (7x – 54)

OpenStudy (anonymous):

ohh :o

OpenStudy (anonymous):

so we get this: = x + 54 mSTR + mSTU = 180 (because it's a linear pair) x + 54 + 8x = 180

OpenStudy (anonymous):

solve from here: x + 54 + 8x = 180

OpenStudy (anonymous):

I got x=14

OpenStudy (anonymous):



OpenStudy (anonymous):

YAYYYYY ~( ^0^)~

OpenStudy (anonymous):

TELL ME WHAT YOU GET MISS CONNECTIONS ACADEMY

OpenStudy (anonymous):

ASDFGHJKJHGFDS how did you know I go to connection ACADEMY

OpenStudy (anonymous):

cause I go there to, I even had that question xD

OpenStudy (anonymous):

And I'm in 9th grade :D